OpenStudy (liizzyliizz):

lim x^3-8/x-2 as x ->2 i know this is easy. but by hand im not getting 12 as the answer. (Something with my factoring i bet) but im tired and i cant think straight help!

OpenStudy (anonymous):

Use L'Hopital's rule..... Or break x^3-8 as (x-2)(x^2-2x+4)

OpenStudy (anonymous):

factor the numerator as the difference of two cubes and cancel, as notsobright wrote. you probably have not made it to l'hopital yet

OpenStudy (anonymous):

however, that is not how to factor

OpenStudy (anonymous):

no it is \[x^3-8=(x-2)(x^2+2x+4)\]

OpenStudy (liizzyliizz):

Curse my algebra skills ! D:

OpenStudy (anonymous):

plus two, not minus two

OpenStudy (anonymous):

oh sorry its +2x not -2x

OpenStudy (anonymous):

then replace x by 2, get 4+4+4=12

OpenStudy (liizzyliizz):

*cry* lol these small fails will be the death of me.

OpenStudy (liizzyliizz):

I know how to do Limits, but the algebra part was driving me insane O.o

OpenStudy (anonymous):

soon you will be able to do this in your head. you will say "the derivative of \[x^3-8\] is \[3x^2\] and \[3\times 2^2=12\]

OpenStudy (anonymous):

yeah the algebra is always the tricky part, idea is easy part

OpenStudy (liizzyliizz):

yeah I understand the calculus part and the concepts. its the factoring that i mess up in, and those negative signs get to me all the time :(